George would like to buy a hi-fi consting R3999. She pays a deposit of 15% and 24 payments of R170 if no insurance premium is added. Calculate the amount of interest in total she would be paying
1
Expert's answer
2020-09-09T17:50:46-0400

"A=P(1+in)"


A="24\times 170= 4080"


"4080=P(1+(0.15\times2))"


"P=\frac {4080}{1.30}"


"P= R3138.46"


Interest= "R3999- R3138.46 = R860.54"


"Interest= R860.54."